The length of a rectangle is six times its width.
If the area of the rectangle is 486yd^2 , find its perimeter

Answer by stanbon(75887)

The length of a rectangle is six times its width.
If the area of the rectangle is 486yd^2 , find its perimeter
-------------------------------
Equations:
L = 6W
---
Area = LW
486 = 6W*W
6W^2 = 486
w^2 = 81
Positive solution:
W = 9 yd
---
L = 6W = 6*9 = 54 yd
---
Perimeter = 2(L+W)
P = 2(9+54)
P = 2(63)
P = 126 yd.
==============
Cheers,
Stan H.